Bank customers arrive randomly on week days afternoon at an average of 3.2 customers every 4 minutes. What is the probability of having more than 7 customers in a 4 minute interval on a week day afternoon?


1
Expert's answer
2021-01-29T02:48:17-0500

"P(x) =\frac{ λ^x \times e^{- λ}} { x!} \\

λ = 3.2 \\

x = 7 \\

P(X>7) =1-P(X=0)-P(X=1)-P(X=2)-P(X=3)-P(X=4)-P(X=5)-P(X=6)-P(X=7)"

Poisson distribution was used.